there are 850 douglas fur and ponderosa pine trees in a section of the forest bought by hoodoo logging company. The company paid an average of $300.00 for each douglas fur and $225.00 for each ponderosa pine. If the company paid $217,500.00 for the trees , how many of each kind did the company buy?
Answer by stanbon(48567)

Let number of fur trees be "x".
Then number of pine trees is "850-x"
Value of fur trees is (300)x dollars
Value of pine trees is (225)(850-x)= 191250-225x dollars
EQUATION:
Value of fur + Value of pine = 217500 dollars
300x + 191250-225x=217500
75x = 26250
x = 350 (number of fur trees)
850-x = 500 (number of pine trees)
Cheers,
Stan H.